OpenStudy (anonymous):

Need Help! Functions, : Use any method to find the point of intersection of f(x)=3x-5 and g(x)=-4x+9

OpenStudy (helder_edwin):

just solve \(f(x)=g(x)\) for \(x\).

OpenStudy (anonymous):

How do you get that? And how does that get you the point of intersection? Would it work if I put it out in a table?

OpenStudy (helder_edwin):

intersection means that the two functions have to meet at some point. that is, they have to be equal for some x.

OpenStudy (anonymous):

So if I put it in a table it at some point they both would have the same x and y intercept?

OpenStudy (helder_edwin):

yes. eventually u would find an \(x_0\) for which \(f(x_0)=g(x_0)\)

OpenStudy (anonymous):

Having trouble making a table with the information,

OpenStudy (anonymous):

Where do I go from there? Sorry lol. Having a brainfart.

hero (hero):

Hint: To find the intersection f(x) = g(x), set 3x-5 = -4x+9

hero (hero):

Solve to find the (x,y) point. In this case, y = f(x)=g(x)

OpenStudy (anonymous):

I'm not sure how to find the y point, or would it be (-5) and 9? and then just insert whatever I solve for x into the x of (x,y)?